# Computational Modelling of Thixotropic Multiphase Fluids

### v1

@@ Appendix B Droplet size distribution
We performe numerical simulations using the multiphase SDPD methodology proposed in this work for the analysis of LLPS phenomena and examined four main properties: (i) protein phase volume fraction $\Phi_{d}$, (ii) capillary velocity $Ca_{v}=\sigma_{o}/\eta_{s}$, (iii) constitutive constant ${1}/{1+\alpha}$, and (iv) thermal energy $k_{B}T$. For each property we construct histograms and estimate the probability density using the KDE function. In addition, we calculate the arithmetic mean, the median and the Weight Average Molecular Weight. In this appendix we present the histograms corresponding to each property, accompanied by a representative snapshot of the last simulated time. The results for protein phase volume fraction $\Phi_{d}=[5\%,35\%]$, $Ca_{v}=[0.1,2]$, ${1}/{1+\alpha}=[0.5,0.01]$ and $k_{B}T=[0.05,0.1]$ are shown in Figure 13.
@@
(c) Capillary velocity $Ca=0.1$
(d) Capillary velocity $Ca=2$
@@
Figure 13: Cluster histograms for the comparison of (a) protein phase concentration, (b) capillary velocity, (c) constitutive constant $\alpha$ and (d) thermal fluctuations.

### v2

@@ Appendix B Droplet size distribution
We perform numerical simulations using the multiphase SDPD methodology proposed in this work for the analysis of LLPS phenomena and examined four main properties: (i) protein phase volume fraction $\Phi_{d}$, (ii) variation of the ratio between the thixotropic characteristic time and interfacial tension time scale $\lambda_{\text{thix}}/\tau_{\sigma}$, (iii) constitutive constant ${1}/{(1+\alpha)}$, and (iv) thermal energy $k_{B}T$. For each property we construct histograms and estimate the probability density using the KDE function. In addition, we compute the Probability Density Function (PDF) of the droplets size, and introduce a set of relevant metrics, such as the average molecular weight (AM), the median (MED), the weighted-average molecular weight (WA), and polydispersity index (PDI). These metrics are defined as follows
where $N_{\text{drop}}=\sum N_{i}$ and $M_{i}$ is the size of the $i$-th droplet. In this appendix we present the histograms corresponding to each property, accompanied by a representative snapshot of the last simulated time. The results for protein phase volume fraction $\Phi_{d}=[5\%,35\%]$, $\lambda_{\text{thix}}/\tau_{\sigma}=\{0.1,0.5,1,2\}$, ${1}/{(1+\alpha)}=[0.5,0.01]$ and $k_{B}T=[0.05,0.1]$ are shown in Figure 14.
In Figure 13, we compile the different Kernel Density Estimation (KDE) curve obtained for the different volume fractions. An snapshot of the final state for two concentration is included. In B (Figure 14(a) and Figure 14(b)), we provide an example of droplet sizes histograms for $\Phi_{d}=5\%$ and $\Phi_{d}=35\%$, illustrating the central tendency statistics. The results indicate that at the lowest volume fractions ($\Phi_{d}=5\%$) investigated, interfacial tension effects are not sufficient to drive the formation of well-defined droplets. In contrast, at higher volume fractions ($\Phi_{d}=35\%$), the protein phase exhibits a tendency to form regular circular droplets, driven by interfacial tension phenomena. Overall, at lower proteins content lead to greater heterogeneity (broader droplet size), whereas as the volume increases, the droplet size distribution tend exhibit lower dispersions.
Figure 13: KDE comparison for four different protein concentration including the PID value. Inside snapshot illustrate the final configuration for two different system, where droplets are depicted in different colors for clarity.
@@
(c) Ratio $\lambda_{\text{thix}}/\tau_{\sigma}=0.1$
(d) Ratio $\lambda_{\text{thix}}/\tau_{\sigma}=2$
@@
Figure 14: Cluster histograms for the comparison of (a) protein phase concentration, (b) variation of the ratio between the thixotropic characteristic time and interfacial tension time scale $\lambda_{\text{thix}}/\tau_{\sigma}$, (c) constitutive constant ${1}/{(1+\alpha)}$, and (d) thermal energy $k_{B}T$.
Appendix C Effective viscosity in suspensions
The behavior of rigid particles suspended in a fluid was studied by Einstein einstein1906 , where the expression $\eta_{eff}=\eta(1+B\phi)$ was derived. This indicates that for concentrations less than 1 ($\phi<<1$) the effective viscosity of the particle suspension is directly proportional to the viscosity of the fluid and the volume fraction of rigid particles, with a constant value $B=2.5$. For deformable particles, Taylor taylor1932viscosity proposed:
where $\eta^{\prime}$ is the viscosity of the dispersed phase and $D$ is the system dimension. When $\eta^{\prime}=\eta$, this simplifies to $\eta_{eff}=\eta(1+1.75\phi)$. This model assumes moderate droplet deformability, i.e., low Capillary numbers. We conduct simulations in a square channel of size $L/\Delta x=50$ under simple shear flow $\dot{\gamma}=0.0166$, and $k_{B}T\to 0$. The continuous phase is Newtonian with $\rho=1$, $\eta=15$ and surface tension $\sigma_{0}=2$. Droplets have radius $r_{0}=0.8$ and viscosities $\eta^{\prime}=\{15\text{ and }30\}$ covering two viscosity ratios of $\lambda=\{1\text{ and }2\}$. This leads to $Re=1.5\times 10^{-3}$ and $Ca=0.1$. We perform a total of eleven simulation with volume fractions in the range $0<\phi<0.3$ by adjusting the number of droplets. The effective viscosity is calculated by direct measurement of the force on the wall. Figure 15 shows the results, including snapshots for $\phi=[0.02,0.14,0.29]$, illustrating droplets deformation. The model captures the increasing trend in viscosity, followed by a slight drop for volume fraction exceeding 0.2, consistent with prior observations caserta2012vorticity , de2019effect , rosti2021shear , rosti2019numerical .
Figure 15: Emulsion validation using Taylor‘s expression for two viscosity ratios $\lambda$